OpenStudy (anonymous):

Need help with a question!!! A number is 5 more than 4 times another. The greater number decreased by 2/3, the smaller equals 20. Find the numbers.

OpenStudy (isaiah.feynman):

@whpalmer4 got a better idea?

OpenStudy (anonymous):

plzz help me.

OpenStudy (isaiah.feynman):

 FIRST STATEMENT.

OpenStudy (anonymous):

ok

OpenStudy (whpalmer4):

is that the greater number - (2/3)* the smaller number = 20?

OpenStudy (isaiah.feynman):

We need another equation. Which seems difficult to set up.

OpenStudy (whpalmer4):

If the problem is as I suggested, then here's how to solve it: let y be the smaller number, and x the larger: "5 more than 4 times another" \[x = 4y + 5\] "the greater number, decreased by 2/3 of the smaller, equals 20" \[x-\frac{2}{3}y = 20\] Now substitute the first equation into the second and solve for \(y\)

OpenStudy (isaiah.feynman):

Ok finally help. So solve the system of equations!

OpenStudy (anonymous):

Thank u so much guys!!!(:

OpenStudy (whpalmer4):

so, do you have an answer for me to check?

OpenStudy (anonymous):

would it be 4 1/2???

OpenStudy (isaiah.feynman):

You need two numbers, because of two unknowns.

OpenStudy (whpalmer4):

yes, 4 1/2 is correct for one of the numbers, what is the other?

OpenStudy (anonymous):

would it be 23

OpenStudy (whpalmer4):

correct again!

OpenStudy (anonymous):

Thanks for all the help and the medal!!!

OpenStudy (whpalmer4):

2/3 of 9/2 = 18/6 = 3, so 23-3 = 20. and 4*9/2 + 5 = 36/2 + 5 = 18+5 = 23, so the solution checks out.